# Zeroth-order Informed Fine-Tuning for Diffusion Model: A Recursive Likelihood Ratio Optimizer

### v2

@@ A.1.1 Prompts
A.1.2 Reward Models and Evaluation Metrics
@@ A.1.3 Baselines
VADER. VADER (Video Diffusion via Reward Gradients) fine-tunes video diffusion models by backpropagating gradients from pre-trained reward models. It enhances sample and computational efficiency, using reward models to assess aesthetics, text-video alignment, and other video-specific tasks. VADER maintains temporal consistency, and generalizes well to unseen prompts, making it an effective tool for adapting video models to complex objectives.
A.1.4 Orthogonal tricks
@@ A.2 Hypeparameters
A.3 Qualitative Results of Text2Image
@@
A.4 Qualitative Results of Text2Video
@@
A.5 Details for Diffusive Chain-of-Thought experiment
Original prompts for the hand task: 1, A realistic open palm facing upward. 2, A picture of a hand facing downward. 3, A hand in a relaxed position. 4, A hand. 5, A photo of a hand.
@@ A.5 Details for Diffusive Chain-of-Thought experiment
A.6 Memory Profile
@@
RLR 22.4 GB 0 22.4 GB
A.7 Proofs
A.7.1 Proof of Proposition 4.1
@@ A.7 Proofs
The interchange of derivative and expectation can be justified by the dominated convergence theorem (Rudin, 1987). Furthermore, the condition of uniform integrability, which is necessary to ensure the safe exchange of limit operations with expectation, can be relaxed by employing a Lipschitz condition (Glasserman, 1990).
By chain rule and $\bm{x}_{0}(z;\theta)=\phi_{1:T}(\bm{x}_{T},z;\theta)=\phi_{1}(\bm{x}_{1},\bm{z}_{1};\theta)$, we have
$\frac{\partial R(\bm{x}_{0}(z;\theta))}{\partial\theta}$ $=\frac{\partial\bm{x}_{0}(z;\theta)}{\partial\theta}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x_{0}}}=\bigg{[}\frac{\partial\phi(\bm{x}_{1},\bm{z}_{1};\theta)}{\partial\theta}+\frac{\partial\phi(\bm{x}_{1},\bm{z}_{1};\theta)}{\partial\bm{x}_{1}}\frac{\partial\bm{x}_{1}(\bm{z}_{2:T};\theta)}{\partial\theta}\bigg{]}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x_{0}}}$
$=\bigg{[}\frac{\partial\phi(\bm{x}_{1},\bm{z}_{1};\theta)}{\partial\theta}+\frac{\partial\phi(\bm{x}_{1},\bm{z}_{1};\theta)}{\partial\bm{x}_{1}}\bigg{[}\frac{\partial\phi(\bm{x}_{2},\bm{z}_{2};\theta)}{\partial\theta}+\frac{\partial\phi(\bm{x}_{2},\bm{z}_{2};\theta)}{\partial\bm{x}_{2}}\frac{\partial\bm{x}_{2}(\bm{z}_{3:T};\theta)}{\partial\theta}\bigg{]}\bigg{]}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x_{0}}}$
=$\dots$= …
$=\bigg{[}\frac{\partial\phi(\bm{x}_{1};\theta)}{\partial\theta}+\sum_{i=2}^{T}\frac{\partial\phi(x_{i};\theta)}{\partial\theta}\prod_{j=1}^{i-1}\frac{\partial\phi(\bm{x}_{j};\theta)}{\partial\bm{x}_{j}}\bigg{]}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x_{0}}}.$
@@
$\nabla_{\theta}\mathbb{E}_{z}[R(\bm{x}_{0}(z;\theta))]=\mathbb{E}_{z}[\nabla_{\theta}R(\bm{x}_{0}(z;\theta))]=\mathbb{E}_{z}\bigg{[}\bigg{[}\frac{\partial\phi(\bm{x}_{1};\theta)}{\partial\theta}+\sum_{i=2}^{T}\frac{\partial\phi(x_{i};\theta)}{\partial\theta}\prod_{j=1}^{i-1}\frac{\partial\phi(\bm{x}_{j};\theta)}{\partial\bm{x}_{j}}\bigg{]}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x_{0}}}\bigg{]},$ (13)
which means the unbiasedness of the FO estimator.
Furthermore, the truncated BP estimator is
$\nabla_{\theta}R(\bm{x}_{0})_{\mathrm{truncated}}=\bigg{[}\frac{\partial\phi_{1}(\bm{x}_{1},\bm{z}_{1};\theta)}{\partial\theta}+\sum_{i=2}^{K}\frac{\partial\phi_{i}(\bm{x}_{i},\bm{z}_{i};\theta)}{\partial\theta}\prod_{j=1}^{i-1}\frac{\partial\phi(\bm{x}_{j},\bm{z}_{j};\theta)}{\partial\bm{x}_{j}}\bigg{]}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x_{0}}}.$ (14)
@@
$\nabla_{\theta}\mathbb{E}[R(\bm{x}_{0})]-\mathbb{E}[\nabla_{\theta}R(\bm{x}_{0})_{\mathrm{truncated}}]=\mathbb{E}_{z_{1:T}}\bigg{[}\sum_{i=T^{\prime}-1}^{T}\frac{\partial\phi(\bm{x}_{i};\theta)}{\partial\theta}\prod_{j=1}^{i-1}\frac{\partial\phi(\bm{x}_{j};\theta)}{\partial\bm{x}_{j}}\bigg{]}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x_{0}}}.$
So we end the proof. ∎
A.7.2 Proof of Proposition 4.2 and additional assumptions
Assumption A.1.
Define $R(z;\theta)$ as $R(\bm{x}_{0}(z;\theta))$. Assume that $R(z;\theta)$ is differentiable with respect to $\theta$ almost surely, $\mathbb{P}(R(\theta)=r)=0$ for every $r\in\mathbb{R}$, and Lipschitz condition holds for every $\theta_{1}$ and $\theta_{2}$:
@@
Assumption A.2.
For any $\bm{x}_{t}$, whose randomness comes from $z$, the density $f(x_{t};\theta)$ is differentiable with respect to $\theta$, and uniform integrability holds:
$\sup_{\theta}\bigg{|}R(x_{t};\theta)\frac{\partial}{\partial\theta}f(x_{t};\theta)\bigg{|}\leq m_{2}(x_{t}),$
where $m_{2}(x_{t})$ is integrable.
Assumption A.3.
@@
Since $R(z;\theta)$ is the reward function and the random variables $\bm{z}_{1:T}$ are Gaussian distributions in our case, it is easy to check the above assumptions are satisfied. By applying the Theorem 2 in (Cui et al., 2020), we can reach the conclusion. ∎
A.7.3 Proof of Theorem 5.1
@@
The RLR estimator contains three parts: FO estimator terms, HO estimator terms, and ZO estimator terms. For simplicity, we can consider there terms with an FO estimator term, an HO estimator term, and an ZO estimator term. Substituting the specific form of the iteration process, we have
$\bm{x}_{0}=\varphi(\bm{x}_{1},\bm{z}_{1};\theta),\quad\bm{x}_{1}=\varphi(\bm{x}_{2};\theta+\bm{z}_{2}),\quad\bm{x}_{2}=\varphi(\bm{x}_{3};\theta)+\bm{z}_{3},$
where $\bm{x}_{1}=\varphi(\bm{x}_{2};\theta+\bm{z}_{2})$ is an HO estimator term and $\bm{x}_{2}=\varphi(\bm{x}_{3};\theta)+\bm{z}_{3}$ is an ZO estimator term.
Then we have
$\nabla_{\theta}\mathbb{E}_{\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}}[R(\bm{x}_{0})]$ $=\nabla_{\theta}\mathbb{E}_{\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}}[R(\varphi(\bm{x}_{1},\bm{z}_{1};\theta))]=\nabla_{\theta}\mathbb{E}_{\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}}[R(\varphi(\varphi(\bm{x}_{2};\theta+\bm{z}_{2}),\bm{z}_{1};\theta))]$ (15)
$=\nabla_{\theta}\mathbb{E}_{\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}}[R(\varphi(\varphi(\varphi(\bm{x}_{3};\theta)+\bm{z}_{3};\theta+\bm{z}_{2}),\bm{z}_{1};\theta))].$
Here $\theta$ appears three times, once in each $\varphi$. When taking derivative with respect to $\theta$, we need to alternate between the chain rule and the compound function rule. First, we apply chain rule and take derivative with respect $\theta$ in the first layer of $\varphi$. The interchange of derivative and expectation can also be justified by the dominated convergence theorem (Rudin, 1987):
@@
$\nabla_{\theta}\mathbb{E}_{\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}}[R(\bm{x}_{0})]=\mathbb{E}_{\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}}\bigg{[}\frac{\partial\varphi(\bm{x}_{1};\theta)}{\partial\theta}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x_{0}}}\bigg{]}+\nabla_{\theta}\mathbb{E}_{\bm{z}_{2},\bm{z}_{3}}\bigg{[}R(\varphi(\varphi(\bm{x}_{2};\theta+\bm{z}_{2})))\bigg{]},$ (16)
where the first term is the first-order term. Then we continue to decompose the second term:
$\nabla_{\theta}\mathbb{E}_{\bm{z}_{2},\bm{z}_{3}}\bigg{[}R(\varphi(\varphi(\bm{x}_{2};\theta+\bm{z}_{2})))\bigg{]}$ $=\mathbb{E}_{\bm{z}_{2}+\theta\sim\mathcal{N}(\theta,\sigma^{2}I),\bm{z}_{3}}\bigg{[}R(\bm{x}_{0})\nabla_{\theta}\ln f(\bm{z}_{2}+\theta)\bigg{]}$ (17)
$\ \quad+\mathbb{E}_{\bm{z}_{2},\bm{z}_{3}}\bigg{[}\frac{\partial}{\partial\bm{x}_{2}}R(\varphi(\varphi(\bm{x}_{2};\theta+\bm{z}_{2})))\frac{\partial\bm{x}_{2}}{\partial\theta}\bigg{]}$
$=\mathbb{E}_{\bm{z}_{2},\bm{z}_{3}}\bigg{[}R(\bm{x}_{0})\nabla_{\theta}\ln f(\bm{z}_{2})\bigg{]}+\nabla_{\theta}\mathbb{E}_{\bm{z}_{3}}\bigg{[}R(\varphi(\varphi(\varphi(\bm{x}_{3};\theta)+\bm{z}_{3})))\bigg{]}$
$=\mathbb{E}_{\bm{z}_{2},\bm{z}_{3}}\bigg{[}R(\bm{x}_{0})\nabla_{\theta}\ln f(\bm{z}_{2})\bigg{]}+\nabla_{\theta}\mathbb{E}_{x_{2}\sim\mathcal{N}(\varphi(\bm{x}_{3};\theta),\sigma^{2}I)}\bigg{[}R(\varphi(\varphi(\bm{x}_{2})))\bigg{]}$
$=\mathbb{E}_{\bm{z}_{2},\bm{z}_{3}}\bigg{[}R(\bm{x}_{0})\nabla_{\theta}\ln f(\bm{z}_{2})\bigg{]}+\mathbb{E}_{\bm{z}_{3}}\bigg{[}R(\bm{x}_{0})D^{\top}_{\theta}\varphi(\bm{x}_{3};\theta)\nabla_{\theta}\ln f(\bm{z}_{3})\bigg{]}.$
Here, the first equation and the fourth equation come from the unbiasedness of LR estimator. That is to say,
$\nabla_{\theta}\mathbb{E}_{z\sim f(z;\theta)}[R(z)]=\mathbb{E}_{z\sim f(z;\theta)}[R(z)\nabla_{\theta}\ln f(z)].$
Combine Equation (16) and (17), we can get the unbiasedness of RLR estimator with three terms:
$\nabla_{\theta}\mathbb{E}_{\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}}\bigg{[}R(\varphi(\varphi(\varphi(\bm{x}_{3};\theta)+\bm{z}_{3};\theta+\bm{z}_{2}),\bm{z}_{1};\theta))\bigg{]}$
$=\mathbb{E}_{\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}}\bigg{[}\frac{\partial\varphi(\bm{x}_{1};\theta)}{\partial\theta}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x_{0}}}\bigg{]}+\mathbb{E}_{\bm{z}_{2},\bm{z}_{3}}\bigg{[}R(\bm{x}_{0})\nabla_{\theta}\ln f(\bm{z}_{2})\bigg{]}+\mathbb{E}_{\bm{z}_{3}}\bigg{[}R(\bm{x}_{0})D^{\top}_{\theta}\varphi(\bm{x}_{3};\theta)\nabla_{\theta}\ln f(\bm{z}_{3})\bigg{]}.$

### v3

@@ B.2 Prompts
B.3 Reward Models and Evaluation Metrics
@@ B.4 Baselines
VADER. VADER (Video Diffusion via Reward Gradients) fine-tunes video diffusion models by backpropagating gradients from pre-trained reward models. It enhances sample and computational efficiency, using reward models to assess aesthetics, text-video alignment, and other video-specific tasks. VADER maintains temporal consistency and generalizes well to unseen prompts, making it an effective tool for adapting video models to complex objectives.
B.5 Orthogonal tricks
@@ Appendix C Hypeparameters
Appendix D Qualitative Results of Text2Image
@@
Appendix E Qualitative Results of Text2Video
@@
Appendix F Details for Diffusive Chain-of-Thought experiment
Original prompts for the hand task: (1) A realistic open palm facing upward. (2) A picture of a hand facing downward. (3) A hand in a relaxed position. (4) A hand. (5) A photo of a hand. The generated DCoT prompts are in Figure 10. We show the instructions to generate the multi-scale DCoT prompts in Figure 11.
@@ Appendix F Details for Diffusive Chain-of-Thought experiment
Figure 10: Qualitative examples for DCoT prompts
Figure 11: Instruction to generate DCoT prompts.
Appendix G Memory Profile, Time Complexity, and Different Diffusion Solvers
@@
RLR 22.4 GB 0 GB 22.4 GB
We report the per-step time cost for the truncated BP method, AlignProp, and compare the three methods in Table 7. As shown in the following table, AlignProp has the highest per-step cost, while RL is the fastest. When considering the total time to reach the same performance level (AES = 5.4), the advantage of RLR becomes more evident. RL requires significantly more steps to converge due to high-variance gradients, whereas RLR achieves the target in fewer steps thanks to its low-variance property, resulting in faster convergence than both RL and AlignProp.
Table 7: Time complexity of Text2Image experiments on SD 1.4
Methods RLR RL Alignprop
Time(min/step) 1.61 0.82 2.85
Time(to same score) 121 (75 steps) 492 (600 steps) 285 (100 steps)
We use RLR to train SD1.4 with the DPMSolverMultistepScheduler to verify its effectiveness on different samplers. We set the inference steps to 20 and the solver order to 2. The results in the Table 8 show that RLR maintains robust performance across both samplers. Diffusion-based generation methods such as DDIM are deterministic when $\eta=0$, and only become stochastic when $\eta>0$. In our experiments, when we require the HO estimator at a given step, we explicitly set $\eta=0.1$ for the step to perform HO. For all other steps, we retain the deterministic ODE solver structure by setting $\eta=0$. This approach allows us to flexibly combine deterministic and stochastic transitions within the same framework, ensuring the correctness of the HO estimator where needed. DPM-Solver supports both deterministic and stochastic sampling modes. Stochasticity can be selectively enabled at the specific steps where the HO estimator is used, with all other steps remaining deterministic.
Table 8: Ablation of different diffusion solver in Text2Image experiments on SD 1.4
Methods PickScore HPSv2 AES ImageReward
DDIM 20.14 28.57 6.53 75.65
DPMSolver 20.21 28.55 6.56 75.70
Appendix H Ablations on the Sub-chain Length $h$
Table 9: Comparison of methods on HPSv2 and ImageReward with memory(GB) and time cost(minute per step).
Method HPSv2 ImageReward Memory Wall clock time
ZO 22.31 40.82 11.8 0.79
DDPO (RL) 22.79 52.06 12.4 0.82
RLR($h=0$) 26.70 63.85 12.7 0.90
RLR($h=1$) 28.02 69.85 18.8 1.15
RLR($h=2$) 29.22 76.55 22.4 1.61
RLR($h=3$) 29.36 76.62 32.6 5.65
RLR($h=4$) 29.55 76.70 45.8 9.23
We provide ablation over different sub-chain lengths $h$ and different estimators in the following two tables. We report the reward score from two models, memory, and wall clock time. As the $h$ increases, the performance increases due to the reduced variance. Increasing $h$ would also increase the computation overhead, including memory and time. We find that when $h$ is larger than 2, the performance gain is marginal compared to the increasing computation burden (high memory requirement and long wall clock time). When the $h$ is smaller than 2, high variance would lead to performance degradation. Therefore, we choose $h=2$ as the best hyperparameter in the main experiment.
Appendix I More Results on Vbench
VBench Evaluation Dimensions.
VBench evaluates video generation models across 16 disentangled dimensions, categorized into Quality and Semantic groups.
The Quality category consists of 7 metrics: Subject Consistency and Background Consistency measure identity and scene consistency across frames using feature similarity; Motion Smoothness evaluates physically plausible motion via a motion prior model; Dynamic Degree quantifies motion magnitude to penalize overly static videos; Temporal Flickering measures high-frequency temporal instability; Aesthetic Quality reflects perceptual appeal using an aesthetic predictor; and Imaging Quality assesses distortions like blur or noise using an image quality model.
The Semantic category includes 9 metrics: Object Class and Multiple Objects test object presence and compositionality; Human Action verifies accurate motion execution; Color, Spatial Relationship, and Scene check fidelity to prompt-specified attributes and layouts; Appearance Style and Temporal Style assess stylistic alignment in space and time; and Overall Consistency captures general text-video correspondence. Each dimension has tailored prompts and automatic evaluation pipelines, ensuring fine-grained, human-aligned assessment.
According to the VBench protocol, the Total Score(TS) is computed as a weighted sum of the Quality Score(QS) and Semantic Score(SS), following the formula: $\text{TS}=0.8\cdot\text{QS}+0.2\cdot\text{SS}.$ In our evaluation, the proposed RLR method achieves the highest scores across all three levels: QS, SS, and the final TS. Notably, most metrics contributing heavily to QS are reported in the main text due to their strong correlation with perceptual quality and their larger weights in the TS calculation. For completeness, we present all 16 VBench metrics in Table 10, where RLR consistently outperforms existing baselines.
Table 10: Automatic evaluation on VBench.
(a) Quality dimensions and total score.
Method Subject Background Motion Dynamic Aesthetic Imaging Temporal Quality Total
Consistency Consistency Smoothness Degree Quality Quality Flickering Score Score
ModelScopeT2V 89.97 89.87 95.79 66.39 52.06 58.57 98.28 78.05 75.75
Open-Sora 92.09 97.39 95.61 48.61 57.76 61.51 98.41 78.82 75.91
Pika 96.76 98.95 99.51 37.22 63.15 62.33 99.77 82.68 80.40
Gen-2 97.61 97.61 99.58 18.89 66.96 67.42 99.56 82.47 80.58
T2V-Turbo 96.28 97.02 97.34 49.17 63.04 72.49 97.48 82.57 81.01
DDPO 95.53 96.63 96.92 58.29 59.23 66.84 97.63 81.43 79.84
VADER 95.79 96.71 97.06 66.94 66.04 69.93 97.62 83.75 81.84
RLR 97.64 97.19 98.05 70.69 66.15 71.08 97.70 85.20 83.14
(b) Semantic dimensions.
Method Object Multiple Human Color Spatial Scene Appearance Temporal Overall Semantic
Class Objects Action Relationship Style Style Consistency Score
ModelScopeT2V 82.25 38.98 92.40 81.72 33.68 39.26 23.39 25.37 25.67 66.54
Open-Sora 74.98 33.64 85.00 78.15 43.95 37.33 21.58 25.46 26.18 64.28
Pika 87.45 46.69 88.00 85.31 65.65 44.80 21.89 24.44 25.47 71.26
Gen-2 90.92 55.47 89.20 89.49 66.91 48.91 19.34 24.12 26.17 73.03
T2V-Turbo 93.96 54.65 95.20 89.90 38.67 55.58 24.42 25.51 28.16 74.76
DDPO 91.36 45.67 94.81 90.22 37.54 55.37 25.27 25.04 28.03 73.47
VADER 91.89 48.67 95.12 91.56 39.78 54.39 25.14 25.23 28.19 74.21
RLR 92.52 50.22 95.20 92.31 40.33 54.18 25.74 25.35 28.27 74.87
Appendix J Proofs
J.1 Proof of Proposition 6.1
@@ Appendix I More Results on Vbench
We assume: $R:\mathbb{R}^{d}\to\mathbb{R}$ and $\phi:\mathbb{R}^{d}\times\mathbb{R}^{m}\times\Theta\to\mathbb{R}^{d}$ are continuously differentiable, and their gradients with respect to the targeted arguments are uniformly bounded. That is, there exist finite constants $M_{R}>0$ and $M_{\phi}>0$ such that $\sup_{x}|\nabla R(x)|\leq M_{R}$, $\sup_{(x_{t},z_{t},\theta)}|\nabla_{(x_{t},\theta)}\phi_{t}(x_{t},z_{t},\theta)|\leq M_{\phi}$. Under these assumptions, for the composite function $R(\phi_{1:T}(x_{T},z_{1:T},\theta))$, the partial derivative with respect to $\theta$ is also uniformly bounded by the chain rule (as a product of bounded terms). Therefore, the integrability condition required by the Dominated Convergence Theorem (Rudin, 1987) is satisfied, which justifies interchanging the gradient and the expectation (Glasserman, 1990).
By chain rule and $x_{0}=\phi_{1:T}(x_{T},z_{1:T};\theta)=\phi_{1}(x_{1},z_{1};\theta)$, we have
$\frac{\partial R(x_{0})}{\partial\theta}$ $=\frac{\partial x_{0}}{\partial\theta}^{\top}\frac{dR(x_{0})}{dx_{0}}=\bigg[\frac{\partial\phi_{1}(x_{1},z_{1};\theta)}{\partial\theta}+\frac{\partial x_{0}}{\partial x_{1}}\frac{\partial x_{1}}{\partial\theta}\bigg]^{\top}\frac{dR(x_{0})}{dx_{0}}$
$=\bigg[\frac{\partial\phi_{1}(x_{1},z_{1};\theta)}{\partial\theta}+\frac{\partial x_{0}}{\partial x_{1}}\bigg[\frac{\partial\phi_{2}(x_{2},z_{2};\theta)}{\partial\theta}+\frac{\partial x_{1}}{\partial x_{2}}\frac{\partial x_{2}}{\partial\theta}\bigg]\bigg]^{\top}\frac{dR(x_{0})}{dx_{0}}$
$=\dots$
$=\bigg[\frac{\partial\phi_{1}(x_{1},z_{1};\theta)}{\partial\theta}+\sum_{i=2}^{T}\frac{\partial\phi_{i}(x_{i},z_{i};\theta)}{\partial\theta}\prod_{j=1}^{i-1}\frac{\partial x_{j-1}}{\partial x_{j}}\bigg]^{\top}\frac{dR(x_{0})}{dx_{0}}.$
@@
$\nabla_{\theta}\mathbb{E}_{z}[R(x_{0})]$ $=\mathbb{E}_{z_{1:T}}[\nabla_{\theta}R(\phi_{1}(x_{1},z_{1};\theta))]$
$=\mathbb{E}_{z_{1:T}}\bigg[\bigg[\frac{\partial\phi(x_{1},z_{1};\theta)}{\partial\theta}+\sum_{i=2}^{T}\frac{\partial\phi_{i}(x_{i},z_{i};\theta)}{\partial\theta}\prod_{j=1}^{i-1}\frac{\partial x_{j-1}}{\partial x_{j}}\bigg]^{\top}\frac{dR(x_{0})}{dx_{0}}\bigg],$
which means the unbiasedness of the FO estimator. Furthermore, the truncated BP estimator is
$\nabla_{\theta}R(\bm{x}_{0})_{\mathrm{truncated}}=\bigg[\frac{\partial\phi_{1}(x_{1},z_{1};\theta)}{\partial\theta}+\sum_{i=2}^{T^{\prime}}\frac{\partial\phi_{i}(x_{i},z_{i};\theta)}{\partial\theta}\prod_{j=1}^{i-1}\frac{\partial x_{j-1}}{\partial x_{j}}\bigg]^{\top}\frac{dR(x_{0})}{dx_{0}}.$ (11)
@@
$\nabla_{\theta}\mathbb{E}[R(x_{0})]-\mathbb{E}[\nabla_{\theta}R(x_{0})_{\mathrm{truncated}}]=\mathbb{E}_{z_{1:T}}\bigg[\sum_{i=T^{\prime}+1}^{T}\frac{\partial\phi(x_{i};\theta)}{\partial\theta}\prod_{j=1}^{i-1}\frac{\partial\phi(x_{j};\theta)}{\partial x_{j}}\bigg]^{\top}\frac{dR(x_{0})}{dx_{0}},$
which completes the proof. ∎
J.2 Proof of Proposition 6.2 and Additional Assumptions
Assumption J.1.
Define $R(z;\theta)$ as $R(\phi_{1:T}(x_{T},z_{1:T};\theta))$. Assume that $R(z;\theta)$ is differentiable with respect to $\theta$ almost surely, $\mathbb{P}(R(\theta)=r)=0$ for every $r\in\mathbb{R}$, and Lipschitz condition holds for every $\theta_{1}$ and $\theta_{2}$:
@@
Assumption J.2.
For any $x_{t}$, whose randomness comes from $z$, the density $f(x_{t};\theta)$ is differentiable with respect to $\theta$, and uniform integrability holds:
$\sup_{\theta}\bigg|R(z;\theta)\frac{\partial}{\partial\theta}f(x_{t};\theta)\bigg|\leq m_{2}(z),$
where $m_{2}(z)$ is integrable.
Assumption J.3.
@@
Since $R(z;\theta)$ is the reward function and the random variables $z_{1:T}$ are Gaussian distributions in our case, it is easy to check that the above assumptions are satisfied. By applying Theorem 2 in Cui et al. (2020), we can reach the conclusion. ∎
J.3 Proof of Theorem 6.3
@@
The RLR estimator contains three parts: FO estimator terms, HO estimator terms, and ZO estimator terms. For simplicity, we can consider the terms with an FO estimator term, an HO estimator term, and a ZO estimator term. Substituting the specific form of the iteration process, we have
$x_{0}=\phi(x_{1},z_{1};\theta),\quad x_{1}=\varphi(x_{2};\theta+z_{2}),\quad x_{2}=\varphi(x_{3};\theta)+z_{3},$
where $x_{1}=\varphi(x_{2};\theta+z_{2})$ is an ZO estimator term and $x_{2}=\varphi(x_{3};\theta)+z_{3}$ is an HO estimator term. We define $z_{2}\sim f_{\text{ZO}}(z)$ and $z_{3}\sim f_{\text{HO}}(z)$ to indicate the noise distribution associated with ZO and HO estimators. The overall gradient can be written as:
$\nabla_{\theta}\mathbb{E}_{z_{1:3}}[R(x_{0})]=\sum_{i=1}^{3}\nabla_{\theta_{i}}\mathbb{E}_{z_{1:3}}[R(\phi(\varphi(\varphi(x_{3};\theta_{3})+z_{3};\theta_{2}+z_{2}),z_{1};\theta_{1}))]\bigg|_{\theta_{1}=\theta_{2}=\theta_{3}=\theta}$ (12)
For the FO term, $\phi(x_{1},z_{1};\theta)$, the gradient is derived by applying the chain rule:
$\nabla_{\theta_{1}}\mathbb{E}_{z_{1:3}}$ $[R(\phi(\varphi(\varphi(x_{3};\theta)+z_{3};\theta+z_{2}),z_{1};\theta_{1}))]\bigg|_{\theta_{1}=\theta}$
$=\nabla_{\theta_{1}}\mathbb{E}_{z_{1:3}}[R(\phi(x_{1},z_{1};\theta_{1})]\bigg|_{\theta_{1}=\theta}=\mathbb{E}_{z_{1:3}}\bigg[D_{\bar{\theta}}^{\top}{\phi(x_{1},z_{1};\bar{\theta})}\frac{dR(x_{0})}{dx_{0}}\bigg]\bigg|_{\bar{\theta}=\theta}$
To derive the gradient of the ZO and HO terms, the key idea is to push the parameter of interest into the density function of the corresponding noise via a change of variables, so that the gradient operator only acts on the log-density function. We have the gradient of the ZO term:
$\nabla_{\theta_{2}}\mathbb{E}_{z_{1},z_{2},z_{3}}$ $[R(\phi(\varphi(\varphi(x_{3};\theta)+z_{3};\theta_{2}+z_{2}),z_{1};\theta))]\bigg|_{\theta_{2}=\theta}$
$=\nabla_{\theta_{2}}\mathbb{E}_{z_{1},z_{2},z_{3}}[R(\phi(\varphi(x_{2};\theta_{2}+z_{2}),z_{1};\theta))]\bigg|_{\theta_{2}=\theta}$
$=\nabla_{\theta_{2}}\mathbb{E}_{z_{1},z_{3}}[\mathbb{E}_{z_{2}\sim f_{\text{ZO}}(z)}[R(\phi(\varphi(x_{2};\theta_{2}+z_{2}),z_{1};\theta))|z_{1},z_{3}]]\bigg|_{\theta_{2}=\theta}$
$=\nabla_{\theta_{2}}\mathbb{E}_{z_{1},z_{3}}[\mathbb{E}_{v_{2}\sim f_{\text{ZO}}(v-\theta_{2})}[R(\phi(\varphi(x_{2};v_{2}),z_{1};\theta))|z_{1},z_{3}]]\bigg|_{\theta_{2}=\theta}$
$=\mathbb{E}_{z_{1},z_{3}}[\mathbb{E}_{v_{2}\sim f_{\text{ZO}}(v-\theta_{2})}[R(x_{0})\nabla_{\theta_{2}}\ln f_{\text{ZO}}(v_{2}-\theta_{2})|z_{1},z_{3}]]\bigg|_{\theta_{2}=\theta}$
$=\mathbb{E}_{z_{1},z_{3}}[\mathbb{E}_{z_{2}\sim f_{\text{ZO}}(z)}[-R(x_{0})\nabla_{z}\ln f_{\text{ZO}}(z_{2})|z_{1},z_{3}]]\bigg|_{\theta_{2}=\theta}$
$=\mathbb{E}_{z_{1},z_{2},z_{3}}[-R(x_{0})\nabla_{z}\ln f_{\text{ZO}}(z_{2})],$
[…]
$\nabla_{\theta_{3}}\mathbb{E}_{z_{1},z_{2},z_{3}}$ $[R(\phi(\varphi(\varphi(x_{3};\theta_{3})+z_{3};\theta+z_{2}),z_{1};\theta))]\bigg|_{\theta_{3}=\theta}$
$=\nabla_{\theta_{3}}\mathbb{E}_{z_{1},z_{2}}[\mathbb{E}_{z_{3}\sim f_{\text{HO}}(z)}[R(\phi(\varphi(\varphi(x_{3};\theta_{3})+z_{3};\theta+z_{2}),z_{1};\theta))|z_{1},z_{2}]]\bigg|_{\theta_{3}=\theta}$
$=\nabla_{\theta_{3}}\mathbb{E}_{z_{1},z_{2}}[\mathbb{E}_{v_{3}\sim f_{\text{HO}}(v-\varphi(x_{3};\theta_{3}))}[R(\phi(\varphi(v_{3};\theta+z_{2}),z_{1};\theta))|z_{1},z_{2}]]\bigg|_{\theta_{3}=\theta}$
$=\mathbb{E}_{z_{1},z_{2}}[\mathbb{E}_{v_{3}\sim f_{\text{HO}}(v-\varphi(x_{3};\theta_{3}))}[R(x_{0})\nabla_{\theta_{3}}\ln f_{\text{HO}}(v_{3}-\varphi(x_{3};\theta_{3}))|z_{1},z_{2}]]\bigg|_{\theta_{3}=\theta}$
$=\mathbb{E}_{z_{1},z_{2}}\bigg[\mathbb{E}_{z_{3}\sim f_{\text{HO}}(z)}[-R(x_{0})D_{\theta_{3}}^{\top}\varphi(x_{3};\theta_{3})\nabla_{z}\ln f_{\text{HO}}(z_{3})|z_{1},z_{2}]\bigg]\bigg|_{\theta_{3}=\theta}$
$=\mathbb{E}_{z_{1},z_{2},z_{3}}\big[-R(x_{0})D_{\bar{\theta}}^{\top}\varphi(x_{3};\bar{\theta})\nabla_{z}\ln f_{\text{HO}}(z_{3})\big]\bigg|_{\bar{\theta}=\theta}.$
Finally, we have the unbiased RLR gradient estimator:
@@
$\nabla_{\theta}\mathbb{E}_{z_{1:3}}[R(x_{0})]=\mathbb{E}_{z_{1:3}}\bigg[\frac{\partial\phi(x_{1},z_{1};\bar{\theta})}{\partial\bar{\theta}}^{\top}\frac{dR(x_{0})}{dx_{0}}-R(x_{0})\bigg(\nabla_{z}\ln f_{\text{ZO}}(z_{2})+\frac{\partial\varphi(x_{3};\bar{\theta})}{\partial\bar{\theta}}^{\top}\nabla_{z}\ln f_{\text{HO}}(z_{3})\bigg)\bigg]\bigg|_{\bar{\theta}=\theta}.$